cos^2theta - cos theta = 0
I don't understand how to solve this and how to do the work.

Found 2 solutions by harpazo, ikleyn:
Answer by harpazo(655)

cos^2 θ - cos θ = 0
Factor out cos θ.
cos θ(cos θ - 1) = 0
Set each factor to 0 and solve for θ.
cos θ = 0
θ = pi/2, 3pi/2
cos θ - 1 = 0
cos θ = 1
θ = 2pi

Answer by ikleyn(53765)